OpenStudy (anonymous):

help with logic... 2+3≠5 if and only if 10-1=3^2. True or false?/why?

ganeshie8 (ganeshie8):

break it into two statements, then test the truth value of each statement

ganeshie8 (ganeshie8):

if you get same truth value for both statements, then the truth value of if and only if is True, otherwise False.

ganeshie8 (ganeshie8):

`2+3≠5 if and only if 10-1=3^2` 2+3≠5 True/False ? 10-1=3^2 True/False ?

OpenStudy (anonymous):

so first statement is false, and second is true, and the rule of 'if and only if' means that the entire statement is False?

ganeshie8 (ganeshie8):

You got it !

ganeshie8 (ganeshie8):

for `if and only if` to be true, both statements must have the SAME truth value.

OpenStudy (anonymous):

Oh I get it now thank you so much!!

ganeshie8 (ganeshie8):

np :)